# ChartBlender: An Interactive System for Authoring and Synchronizing Visualization Charts in Video

### v1

@@ 6 Evaluation
We evaluated the ChartBlender and the key techniques through three approaches: (1) a quantitative experiment to assess the precision of camera tracking, (2) a quantitative experiment to assess the precision of object tracking, and (3) domain expert interviews to evaluate the overall usability of the system.
6.1 Camera Tracking Precision
In our system, camera tracking is implemented using RGBD odometry, which estimates frame-by-frame camera motion by leveraging both RGB appearance and depth information. As a result, the quality of depth estimation plays a critical role in the precision of camera tracking. To evaluate this dependency, we compare how different monocular depth models affect the overall pose estimation performance of the RGBD odometry pipeline.
Experimental Setup. To evaluate the impact of depth estimation quality on camera pose recovery, we conducted a comparative experiment using different monocular depth models combined with RGBD odometry. Specifically, we compared Metric3D, which emphasizes metric depth precision, with ZoeDepth, which focuses on generalization and temporal consistency. We refer to the two pipelines as Metric3D + RGBD odometry (our proposed method) and ZoeDepth + RGBD odometry (baseline). We selected a subset of the DriveTrack dataset [43], consisting of 18 video sequences (1,806 frames in total), where camera motion was sufficiently pronounced and affected by dynamic foreground objects. For each method, predicted depth maps were fed into RGBD odometry implemented in Open3D [39], which uses a hybrid geometric-photometric optimization [40] to estimate per-frame camera extrinsics. A fixed intrinsic matrix of the form [1000, 1000, width/2, height/2] was used for all sequences.
Evaluation Metric. To quantitatively evaluate the precision of our camera tracking approach and assess the influence of different depth estimation models, we measured the geometric discrepancy between the estimated camera poses and the ground-truth extrinsics provided by the TAPVid-3D dataset [34]. For each frame, we compared the predicted camera pose $(R_{\text{pred}},t_{\text{pred}})$ with the corresponding ground-truth pose $(R_{\text{gt}},t_{\text{gt}})$, and computed both rotation and translation errors. The rotation error was defined as the angular difference between the predicted and ground-truth orientations. Specifically, we first computed the relative rotation matrix $\Delta R=R_{\text{gt}}^{-1}R_{\text{pred}}$, and then calculated the rotation error as the angle of this rotation:
$\theta=\arccos\left(\frac{\operatorname{trace}(\Delta R)-1}{2}\right),$ (9)
This formulation captures the minimum angular difference between the two rotation matrices in 3D space. The translation error was defined as the Euclidean distance between the predicted and ground-truth camera positions:
[…]
Together, these two metrics quantify both orientation and positional discrepancies in the estimated camera trajectories. To evaluate overall precision and robustness to outliers, we report both the mean and median rotation and translation errors computed across all frames in the dataset.
Results and Analysis. As shown in Table I, Metric3D consistently outperformed ZoeDepth in both rotation and translation precision when paired with RGBD odometry, achieving lower mean and median errors across all frames. Although both methods operate under the same fixed intrinsic setting, Metric3D provided more reliable metric depth, enabling more accurate frame-by-frame camera pose estimation. Another key reason for adopting Metric3D is its robustness in unconstrained environments where intrinsic calibration is unavailable, a common challenge in user-generated or in-the-wild videos. Despite the use of an approximate intrinsic matrix, the Metric3D + RGBD odometry pipeline maintained strong performance.
@@ 6.1 Camera Tracking Precision
6.2 Object Tracking Precision
We evaluated our system’s object tracking techniques by assessing the precision of the reconstructed 3D motion trajectories for dynamic objects. This reflects how well the system can estimate object movement in physical space.
Experimental Setup. In our system, object tracking is implemented by combining 2D keypoint tracking with monocular depth estimation to recover the 3D trajectories of dynamic objects across time. This process is essential for enabling visualization alignment with moving targets in the scene. The overall tracking precision depends on the quality of both the 2D keypoint trajectories and the estimated depth maps used for lifting to 3D. To evaluate this component, we conducted a systematic comparison of different tracking configurations by varying the underlying 2D tracking and depth estimation models. Specifically, we considered two keypoint tracking models, TAPIR [44] and BootsTAPIR [41], and two depth estimation models, ZoeDepth and Metric3D, resulting in four combinations in total.
We performed this evaluation on a subset of the PStudio dataset [45], consisting of 36 video sequences (10,800 frames in total). These sequences feature clearly observable dynamic objects with prominent motion and minimal camera movement, allowing us to isolate object trajectory estimation from background motion. For each combination, 2D keypoints were first tracked across frames using the selected tracking model. These trajectories were then lifted into 3D space by back-projecting the 2D points with the corresponding depth estimates.
Evaluation Metric. We adopted the Average 3D Point Distance (APD3D) metric, introduced in TAPVid-3D [34], to evaluate the precision of 3D point tracking. This metric quantifies the proportion of predicted 3D points that lie within a specified Euclidean distance threshold of their corresponding ground-truth positions, averaged over time and across tracked points. A higher value reflects better spatial precision and temporal consistency of the reconstructed trajectories.
Formally, the metric is defined as:
$\mathrm{APD}_{3D}\equiv\frac{1}{V}\sum_{i,t}v_{t}^{i}\cdot\mathbf{1}\left(\lVert\hat{P}_{t}^{i}-P_{t}^{i}\rVert<\delta_{3D}(P_{t}^{i})\right),$ (11)
where $\hat{P}_{t}^{i}$ and $P_{t}^{i}$ denote the predicted and ground-truth 3D positions of the $i$-th point at time $t$, $v_{t}^{i}$ is a visibility indicator, and $V$ is the total number of visible points. To quantify performance across different levels of precision, we report APD3D under a fixed set of distance thresholds: $[0.04,0.08,0.16,0.32,0.64]$.
Results and Analysis. As shown in TABLE II, BootsTAPIR consistently outperformed TAPIR across all configurations, confirming its superior capability in 2D temporal keypoint tracking. With regard to depth estimation, our primary objective is not to reconstruct metrically accurate 3D geometry, but to generate perspective-consistent visualizations that preserve spatial coherence between foreground and background elements. In this context, moderate depth errors are acceptable, provided that the overall depth structure remains stable and visually plausible.
@@ 6.2 Object Tracking Precision
BootsTAPIR + Metric3D (ours) 0.3424
TAPIR + Metric3D 0.3416
BootsTAPIR + ZoeDepth 0.3609
TAPIR + ZoeDepth 0.3595
@@ 6.3 Expert Interview
To further evaluate the usability of ChartBlender, a series of interviews with five data journalists (denoted as E1-E5) were performed. They had over 5 years’ video editing experiences on average, and were very familiar with chart-embedded videos. All participants provided informed consent prior to the study.
Procedure. The interviews were performed via an online meeting system. Each interview started by a 10-minute introduction about the system. Participants were first introduced to our system through a brief tutorial delivered by authors, and were given two task:
@@
Task 2 - Open-ended (30 minutes). Participants are asked to create two novel data videos by choosing among 10 short video clips (10” duration on average).
Finally, the experts were asked to finish a questionnaire, followed by an interview after they explored the functionalities of ChartBlender. The whole process lasted for about one hour, and the interviews were recorded for analysis.
@@
Fast Crafting. All experts agreed that our system was highly effective in supporting both camera tracking and object tracking. In particular, E1 commented that “This system can accomplish the work of several hours, or even days, in just two or three minutes.” E2 mentioned that “This system allows me to see the real-time effect of chart edits in the video… I no longer need to repeatedly switch between project files, which used to waste a lot of time.” E5 also noted that “The speed of result generation is so fast that I can quickly create multiple demos to observe where the chart embedding works best in the video.”
@@
A Professional and Elegant System. All the experts were satisfied with the system. Additionally, all of them mentioned that the system was very user-friendly and that they would recommend it to others. Specifically, they appreciated how seamlessly the system integrates chart creation and embedding into videos, accomplishing highly professional tasks (E1, 2, 5). E1 mentioned that “This system can complete highly professional tasks without requiring extensive skills or knowledge.” E5 noted that “the chart can be translated and rotated in 3D, and it integrates seamlessly with the video.” They were also satisfied with the chart creation aspect. E4 noted that “the data mapping feature eliminated the need for selecting data and creating visualizations in 3D software.” E5 commented that “The chart creation process is very user-friendly for me. There are many expressive templates available for me to use … Previously, I had to spend a lot of time on creating visualizations.”
@@ 7 LIMITATIONS AND FUTURE WORK
Despite the encouraging evaluation results, several discussions and limitations were identified during the design and implementation process or mentioned by the experts during interviews. We hope to guide potential future research directions by pointing out these limitations.
@@
Layered Rendering. Although the current system supports depth-aware relationships between charts, Several experts (E1, E3) highlighted the potential of masking functionality that enables visualizations to be occluded by foreground objects in the video. Such a feature would allow charts to appear partially or fully behind scene elements, enhancing realism and spatial coherence.
@@ 8 Conclusion
In this paper, we introduced ChartBlender, an integrated system for authoring, embedding, and synchronizing data visualizations within dynamic video content. By combining computer vision algorithms for camera and object tracking, we enable synchronization between visualizations and video motion. In addition, the system incorporates an interactive workflow to support easy and efficient authoring of videos with visualizations embedded. Quantitative evaluations demonstrate the effectiveness of our tracking algorithms in recovering spatially consistent motion trajectories. Additionally, interviews with domain experts confirm the system’s usability, expressiveness, and potential for professional media production. The evaluation showed the power of ChartBlender system and revealed several limitations of the current system, which will be addressed in the future.

### v2

@@ 6 Evaluation
We evaluated ChartBlender through (1) two quantitative experiments that assess the precision of our two tracking pipelines and (2) interviews with domain experts that assess system usability.
6.1 Evaluation of Two Tracking Pipelines
Our system relies on two tracking pipelines: camera tracking and object tracking. The camera tracking pipeline consists of (A) depth estimation followed by (B1) camera pose estimation. The object tracking pipeline consists of (A) depth estimation followed by (B2) object trajectory estimation. The two pipelines share the same depth estimation stage. However, to the best of our knowledge, there’s no open-source system that provides an end-to-end baseline for motion-synchronized chart embedding. We therefore construct baselines using a component substitution strategy: we replace individual stages with representative models identified in our literature review and evaluate the resulting combinations under identical experimental settings. This strategy allows us to quantify the tracking precision of two pipelines.
6.1.1 Precision of the Camera Tracking Pipeline.
Experimental Setup. The camera tracking pipeline comprises two core stages: depth estimation and camera pose estimation. In ChartBlender, the camera tracking pipeline is implemented as Metric3D + Open3D RGBD odometry [Zhou2018, park2017colored]. To validate the effectiveness of our pipeline, we construct a baseline by substituting the depth estimation component, yielding ZoeDepth + RGBD odometry (baseline). We choose this baseline because Open3D RGB-D odometry is a frequently used and stable open-source camera pose estimator, making it a reliable and fixed module. ZoeDepth [bhat2023zoedepth] is another top-performing depth estimator identified in our literature review. The goal of our experiment is to compare whether our depth estimation model (Metric3D) achieves better performance than ZoeDepth when paired with the same tracking module.
We evaluate on the subset of DriveTrack [koppula2024tapvid] (18 sequences, 1,806 frames) with ground-truth camera poses, where camera motion is sufficiently pronounced, and scenes include dynamic foreground objects. These conditions closely resemble the real-world videos in which our system embeds charts. For each frame, the predicted depth map is fed into RGBD odometry to estimate camera poses.
Evaluation Metric. We report rotation and translation errors between predicted poses $(R_{\text{pred}},t_{\text{pred}})$ and ground truth $(R_{\text{gt}},t_{\text{gt}})$. Rotation error is computed from the relative rotation $\Delta R=R_{\text{gt}}^{-1}R_{\text{pred}}$ as the corresponding angular difference, and translation error is the Euclidean distance $\lVert t_{\text{gt}}-t_{\text{pred}}\rVert$.
Results and Analysis. Table I shows that our pipeline yields lower rotation and translation errors than the baseline (mean and median). Since both variants use the same odometry backend and camera intrinsics, the improvement is attributable to the depth backbone, indicating that Metric3D provides depth estimates that better support downstream pose recovery in our pipeline.
@@ 6.1 Evaluation of Two Tracking Pipelines
6.1.2 Precision of the Object Tracking Pipeline.
Experimental Setup. Object tracking comprises two stages: depth estimation and object trajectory estimation. In ChartBlender, we implement this pipeline as Metric3D + BootsTAP, which is not the only reasonable choice. During our literature review, we also identified strong alternatives for each stage, including ZoeDepth[bhat2023zoedepth] for depth estimation and TAPIR[doersch2023tapir] for object trajectory estimation. To validate that our chosen pipeline remains the best-performing configuration under the same experimental conditions, we evaluate three additional baseline combinations: Metric3D + TAPIR, ZoeDepth + TAPIR, and ZoeDepth + BootsTAP.
We performed this evaluation on a subset of the PStudio dataset [joo2015panoptic], which contains high-fidelity multiview recordings of social interaction scenes. We selected 36 video sequences (10,800 frames total) that include clearly visible dynamic objects undergoing complex 3D motion, while the camera remains mostly stationary. This setup allows us to isolate the quality of object trajectory reconstruction from the effects of camera movement.
Evaluation Metric. We adopted the Average 3D Point Distance (APD3D) metric [koppula2024tapvid], to evaluate the precision of 3D point tracking. APD3D measures the fraction of predicted 3D points that fall within a given Euclidean distance threshold of the corresponding ground-truth points, averaged over time and across tracks. Higher values indicate more accurate and temporally consistent 3D trajectories.
Results and Analysis. As shown in Table II, pipelines using BootsTAP consistently outperform those using TAPIR under the same depth backbone, indicating that BootsTAP provides more reliable 2D keypoint trajectories for downstream 3D reconstruction.
Although ZoeDepth + BootsTAP achieves the highest APD3D score, we adopt Metric3D + BootsTAP in ChartBlenderfor better alignment with our camera tracking pipeline. Because both variants use the same 2D tracker (BootsTAP), the APD3D gap primarily reflects the choice of depth backbone. With regard to depth estimation, our objective is not metrically exact 3D trajectory recovery, but perspective-consistent chart embeddings that preserve spatial coherence between foreground and background elements. Therefore, moderate depth errors are acceptable, provided that the overall depth structure and recovered trajectories remain stable.
@@
Metric3D + BootsTAP (ours) 0.3424
Metric3D + TAPIR 0.3416
ZoeDepth + BootsTAP 0.3609
ZoeDepth + TAPIR 0.3595
@@ 6.2 Expert Interview
Participants. To further evaluate the usability of ChartBlender, we conducted semi-structured interviews with five data journalists (E1–E5; Male=2, Female=3; Age: 32-45). Two were from our collaborating media organization, while the remaining three were recruited from other newsrooms to broaden perspectives and reduce organizational bias. Participants had, on average, over ten years of video-editing experience and were familiar with chart-embedded videos. All participants provided informed consent before the study.
Procedure. The interviews were performed via an online meeting system. Each interview started with a 10-minute introduction about the system. Participants were first introduced to our system through a brief tutorial delivered by the authors and were given two tasks:
@@
Task 2 - Open-ended (30 minutes). Participants are asked to create two novel data videos by choosing among 10 short video clips (10" duration on average).
Measures. After the tasks, experts participated in a semi-structured interview. Each session lasted approximately 60 minutes. We gathered two data sources: (i) The completed videos, (ii) audio recordings of think-aloud protocols and interviews. Transcripts and observation notes were analyzed using a reflexive thematic analysis. Three authors independently performed open coding to generate candidate codes, reconciled discrepancies through discussion, and iteratively refined a shared codebook until reaching consensus on higher-level themes.
@@
Fast Crafting. All experts agreed that our system was highly effective in supporting both camera tracking and object tracking. In particular, E1 commented that “This system can accomplish the work of several hours, or even days, in just two or three minutes.” E2 mentioned that “This system allows me to see the real-time effect of chart edits in the video… I no longer need to repeatedly switch between project files, which used to waste a lot of time." E5 also noted that “The speed of result generation is so fast that I can quickly create multiple demos to observe where the chart embedding works best in the video."
@@
A Professional and Elegant System. All the experts were satisfied with the system. Additionally, all of them mentioned that the system was very user-friendly and that they would recommend it to others. Specifically, they appreciated how seamlessly the system integrates chart creation and embedding into videos, accomplishing highly professional tasks (E1, 2, 5). E1 mentioned that “This system can complete highly professional tasks without requiring extensive skills or knowledge." E5 noted that “the chart can be translated and rotated in 3D, and it integrates seamlessly with the video.” They were also satisfied with the chart creation aspect. E4 noted that “the data mapping feature eliminated the need for selecting data and creating visualizations in 3D software." E5 commented that “The chart creation process is very user-friendly for me. There are many expressive templates available for me to use … Previously, I had to spend a lot of time on creating visualizations."
@@ 7 LIMITATIONS AND FUTURE WORK
While the evaluation results indicate ChartBlenderis promising to help users embed and synchronize charts with video scenes, the system still has several limitations that were found during the implementation or mentioned by the participants during interviews. We hope to guide potential future research directions by pointing out these limitations.
@@
Semantic Motion Analysis. While our system employs robust pixel‑level tracking to ensure precise geometric synchronization, it lacks the ability to handle occlusions effectively (E1,E2,E3). This limitation stems from its reliance on low‑level visual features without integrating higher‑level semantic reasoning. We regard pixel‑level tracking as a fundamental requirement for stable embedding, as it generalizes well across diverse video types without being constrained by predefined object categories. Introducing semantic recognition could effectively address motion‑occlusion challenges, enabling the system to maintain tracking continuity even when visual features are temporarily obscured.
@@ 8 Conclusion
In this paper, we introduced ChartBlender, an integrated system for authoring, embedding, and synchronizing data visualizations within dynamic video content. By combining computer vision algorithms for camera and object tracking, we enable synchronization between visualizations and video motion. In addition, the system incorporates an interactive workflow to support easy and efficient authoring of videos with visualizations embedded. Quantitative evaluations demonstrate the effectiveness of our algorithm pipeline in recovering spatially consistent motion trajectories. Additionally, interviews with domain experts confirm the system’s usability, expressiveness, and potential for professional media production. The evaluation showed the power of ChartBlender system and revealed several limitations of the current system, which will be addressed in the future.
@@ ACKNOWLEDGMENTS
[Uncaptioned image] Yi He received her B.Eng degree from School of Digital Media & Design Arts, Beijing University of Posts and Telecommunications. She is currently working toward her Ph.D. degree as part of the Intelligent Big Data Visualization (iDVx) Lab, Tongji University. Her research interests include data visualization and artificial intelligence.
[Uncaptioned image] Yuqi Liu received her B.Eng degree from the School of Computer Science and Technology, Shanghai University of Computer Engineering and Science. She is currently working toward her Master’s degree as part of the Intelligent Big Data Visualization (iDVx) Lab at Tongji University. Her research interests include data visualization and artificial intelligence.
[Uncaptioned image] Chenpu Li is pursuing his B.Eng dual degree in Visual Communication Design and Artificial Intelligence at the College of Design and Innovation, Tongji University. His cur-rent academic focus spans Human-Computer Interaction (CHI) and AI-Art integration.
[Uncaptioned image] Ruoyan Chen received her B.Eng degree in Industrial Design from University of Science and Technology Beijing. She is currently pursuing her M.S. degree as part of the Intelligent Big Data Visualization (iDV*) Lab, Tongji University. Her research interests include data visualization and agent design.
[Uncaptioned image] Chuer Chen received her MSc degree from the Department of Electrical and Computer Engineering, National University of Singapore in 2021. She is currently working toward her Ph.D. degree as part of the Intelligent Big Data Visualization (iDVx) Lab, Tongji University. Her research interests include information visualization and intelligent design.
[Uncaptioned image] Shengqi Dang received his B.Sc. in Mathematics from Tongji University and is currently pursuing a Ph.D. degree at the Intelligent Big Data Visualization (iDVx) Lab, Tongji University. His research interests include of artificial intelligence and computer graphics.
[Uncaptioned image] Nan Cao received his Ph.D. degree in Computer Science and Engineering from the Hong Kong University of Science and Technology (HKUST), Hong Kong, China in 2012. He is currently a professor at Tongji University and the Assistant Dean of the Tongji College of Design and Innovation. He also directs the Tongji Intelligent Big Data Visualization Lab (iDVx Lab) and conducts interdisciplinary research across multiple fields, including data visualization, human computer interaction, machine learning, and data mining. He was a research staff member at the IBM T.J. Watson Research Center, New York, NY, USA before joining the Tongji faculty in 2016.